OpenStudy (anonymous):

system of equation y=x^2+3 y=x+5

OpenStudy (anonymous):

if anyone could show me the steps?

OpenStudy (anonymous):

We generally use one to solve the other. Start with simplest. y = x + 5 y(-5) = x+ 5 (-5) y- 5 = x therefore substituting in eq .1 we get: y = (y-5)^2 +3 y = (y^2 -10y +25) + 3 y = y^2 -10y + 22 subtract y from both sides to get: 0 = y^2 -9y + 22 solve to get: